Circular No. 6192

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SATELLITES OF SATURN A. S. Bosh, Lowell Observatory; and A. S. Rivkin, Lowell Observa-

tory and Lunar and Planetary Laboratory, report: “We have discovered four probable satellites of Saturn, from twenty-seven 400-s images taken in the 890-nm methane band with the HST Wide Field Planetary Camera 2 during an 11-hr span at the May 22 ring-plane crossing. For each satellite, we give below the orbital radius; the angular distance from Saturn at May 22.48584 UT (negative values indicating that the satellite is west of Saturn); the longitude at the same epoch (measured from the subearth point); the approximate V magnitude at opposition; and the number of images in which the satellite was detected: S/1995 S 1, 137450 ± 200 km, −16′′.4, +120^{o}, 17.2, 15; S/1995 S 2, 139700 ± 300 km, −18′′.0, +69^{o}, 16.3, 12; S/1995 S 3, 141050 ± 100 km, −10′′.5, +33^{o}, 17.5, 14; S/1995 S 4, 146450 ± 450 km, +14′′.2, −44^{o}, 18.3, 11. S/1995 S 1, S 2, and S 3 were identified in each frame in which they were not close to Saturn or a bright satellite. S/1995 S 4 was not identified in seven frames in which it should have been visible, all before ring-plane crossing, when the background noise was higher. Satellite positions were fitted to a circular orbit in the plane of the rings, including terms arising from Saturn’s nonspherical gravity field. Maximum residuals were in all cases < 0′′.5, and for S/1995 S 2 and S 3 were no greater than 0′′.15, which is approximately the centering error. All known satellites are accounted for except Saturn XVIII (Pan), XV (Atlas), and XVI (Prometheus), which may be visible but are still uncertain; the possibility exists that S/1995 S 1 is Atlas and that S/1995 S 2 is Prometheus, although they would be 26^{o} and 21^{o} from their respective predicted positions. We see no satellites outside the rings (the distance of which from Saturn at the times of observation was > 20′′) with opposition V mag equal to or brighter than 19, i.e., that of Saturn XIV (Calypso).”

V1333 AQUILAE S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence

(OHP), communicate: “CCD photometry of V1333 Aql, the optical counterpart of the recurrent x-ray transient Aquila X-1, has been obtained at Haute-Provence with the 1.2-m telescope at regular intervals since Mar. 11. The object was in quiescence (V = 19.2) until July 17, when it started brightening. Quick-look magnitudes on July 25.0 UT yield V = 17.5, typical for these recurring activity cycles. Monitoring will continue at OHP. Observations are urged at radio and x-ray wavelengths.”

1995 July 26 Daniel W. E. Green